OpenStudy (anonymous):

integrate x^2/1+x^2 dx

OpenStudy (anonymous):

\[\int\limits (1+x^2-1)/(1+x^2)\]

OpenStudy (anonymous):

why did u do that?

OpenStudy (anonymous):

\[\int\limits 1+(1/1+x^2) dx\]

OpenStudy (anonymous):

addind n subtracting 1 in numerator, then splitting up the denominator

OpenStudy (anonymous):

i didnt know u doing like magic

OpenStudy (anonymous):

\[x+ \tan ^{-1} x +c\]

OpenStudy (anonymous):

its there logic process?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

\[\int\limits{{ x^2} \over {1+x^2}} dx\]

OpenStudy (anonymous):

yes just divide. you long division and you will get the same thing.

OpenStudy (anonymous):

is this one u want to evaluate?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

we can have a similar expression in the numerator by just adding n subtracting 1

OpenStudy (anonymous):

ok... but i finish that chapter already and i didnt see my professor doing that...

OpenStudy (anonymous):

\[\int\limits {{ 1+x^2-1} \over {1+x^2}} dx\]

OpenStudy (anonymous):

but why u add and subtract 1.

OpenStudy (anonymous):

now split up the denominator

OpenStudy (anonymous):

give reason

OpenStudy (anonymous):

reason 4 what?

OpenStudy (anonymous):

why u adding 1

OpenStudy (anonymous):

because u have 1+x^2 in the denominator

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok u can keep telling me

OpenStudy (anonymous):

so what u do next

OpenStudy (anonymous):

you really can just divide. really. just divide \[x^2+1\] into \[x^2\] and you will get the same thing. that is if they trick annoys you

OpenStudy (anonymous):

yes satellite is right, go for the long division

OpenStudy (anonymous):

so what should i get?

OpenStudy (anonymous):

1- 1/(1+x^2)

OpenStudy (anonymous):

u got it?